OpenStudy (anonymous):

An ostrich is 108 inches tall. If its shadow is 162 inches, and an emu standing next to it casts a 90-inch shadow, how tall is the emu?

OpenStudy (anonymous):

Look up pythagorean theorem

OpenStudy (anonymous):

This is solvable by using proportions: \[\frac{ 108 }{ 162 } = \frac{ x }{ 90 }\]Cross-multiplying: x = (108)(90)/(162) So, just simplify and you have "x" !

OpenStudy (anonymous):

All good now, @Kolby97 ?

OpenStudy (anonymous):

no @tcarroll010

OpenStudy (anonymous):

The ostrich's shadow is a certain number times its height. Same for the emu. Notice that both fractions, once you simplify to x=60, are the same. Here's a picture:

OpenStudy (anonymous):

In this double triangle, height is proportional to the shadow. It's really all about similar triangles as that lower right angle is the exact same for both. And actually all corresponding angles are the same measure.

OpenStudy (anonymous):

Any better now? @Kolby97 ?

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @Kolby97